# Learning a Distributed Hierarchical Locomotion Controller for Embodied Cooperation

### v1

@@ A.4.1 Corridor Crossing
We record the success rates of each Ant at its initial position separately, as illustrated in Fig. A.4.1. We have a higher success rate when the initial position is closer to the center, and a higher failure rate when the initial position is closer to the edge. This is also explainable, as the Ant at the centre is closer to the slit and does not need to adjust velocity.
A.4.2 Ravine Bridging
The bridging and coordination in the ravine is the key to completing the task. We set up three difficulty levels - easy, medium, and hard - based on the distance from the ravine’s edge. The easy level corresponds to a distance of 2 m, the medium level to 4 m, and the hard level to 6 cm. The Tab.  3 shows the success rates of the three difficulty levels with different numbers of Ants. We can observe that it becomes more difficult for any number of Ants to complete the task as the distance increases. Because cooperation on the timing of stopping and bridging becomes more important when moving length increases.

### v2

@@ A.2.2 Corridor Crossing
We record the success rates of each Ant at its initial position separately, as illustrated in Fig. A.3. We have a higher success rate when the initial position is closer to the center, and a higher failure rate when the initial position is closer to the edge. This is also explainable, as the Ant at the centre is closer to the slit and does not need to adjust velocity.
A.3 Comparison to Other MARL Alogithms
We conduct comparison between the proposed distributed HRL method to other state-of-the-art MARL approaches based on the article [19]. Specifically, we compare our fully-decentralized method with MAPPO and MASAC in task Cooperative Transport(4 agents) since the Isaac Sim algorithm is nested within the RLGame and the CTDE algorithms supported by RLGame are only MAPPO and MASAC. Fig. A.3 shows the results. We find that MAPPO converges a little faster than distributed method because it is centrally trained with global observations. However, the CTDE algorithm has worse performance and take much more training time when facing a larger number of agents, which has been demonstrated in Fig. 5 of the scalability experiments part.
A.4 Training pipeline of Ravine Bridging
As we mentioned in Appendix A.2, the Ravine Bridging task is more intricate, which means we have to provide more guidance. And the agent’s behavior is divided into stages, so the training process must also be divided accordingly. Therefore, there is a series of pre-train courses to complete this task. For the agent in the ravine, its upper layer receives observations that include the position of the bridging stone, its own position, and the position of the Ant on the plane closest to the ravine. The Ant on the plane receives observations that includes the position of the stone, its own position, and the x coordinate of the ravine and the finish line. Before training together, we train them separately first to obtain their featured skills. For the agents in the ravine, we randomize the position of the Ant closest to the ravine in the input and give the agent a y-axis direction distance reward of the distance between the stone and the randomize Ant’s position, so that the agent in the ravine has the ability to move the stone to the closest Ant on the plane. For the agents on the plane, we add a reward for the x-axis direction distance to the finish line and a punishment for falling (death), so that it learns the ability to move straight towards the finish line and stop before falling into the ravine. Finally, we train all of them together, and once an Ant crosses the finish line, we give all of them a large sparse reward. For all the reward precise equations mentioned above, we will list them in the following section.
A.5 Reward Details
A.5.1 Lower layer
Target paradigm
Reward Type Formula
Distance to target reward $5\times\left(\frac{1}{1+\sqrt{(\text{agent\_pos}[:,0]-\text{target\_pos}[:,0])^{2}+(\text{agent\_pos}[:,1]-\text{target\_pos}[:,1])^{2}}}\right)$
Alive reward 0.5
Up reward 0.1
[…]
Velocity paradigm
Reward Type Formula
[…]
Alive reward 0.5
Up reward 0.1
[…]
A.5.2 Upper layer
Cooperative Transport
Reward Type Formula
Cylinder-target distance reward $10\times\left(\frac{1}{1+\sqrt{(\text{cylinder\_pos}[:,0]-\text{target\_pos}[:,0])^{2}+(\text{cylinder\_pos}[:,1]-\text{target\_pos}[:,1])^{2}}}\right)$
Agent-cylinder distance reward $5\times\left(\frac{1}{1+\sqrt{(\text{agent\_pos}[:,0]-\text{cylinder\_pos}[:,0])^{2}+(\text{agent\_pos}[:,1]-\text{cylinder\_pos}[:,1])^{2}}}\right)$
Agent-target distance reward $5\times\left(\frac{1}{1+\sqrt{(\text{agent\_pos}[:,0]-\text{target\_pos}[:,0])^{2}+(\text{agent\_pos}[:,1]-\text{target\_pos}[:,1])^{2}}}\right)$
Reach destination reward 50 (all agents)
Cross Corridor
Reward Type Formula
Reach destination reward 50 (all agents)
Agent-target distance reward $5\times\left(\frac{1}{1+\sqrt{(\text{agent\_pos}[:,0]-\text{target\_pos}[:,0])^{2}+(\text{agent\_pos}[:,1]-\text{target\_pos}[:,1])^{2}}}\right)$
Ravine Bridging
Reward/Punishment Type Formula
Ravine pretrain: Distance of stone and nearest agent reward $5\times\left(\frac{1}{1+|\text{stone\_y}-\text{agent\_y}|}\right)$
Plane pretrain: Distance to destination line reward $5\times\left(\frac{1}{1+|\text{destination\_x}-\text{agent\_x}|}\right)$
Plane pretrain: Falling down punishment -20
Integrated training: Pass destination line reward 50 (all agents)